A cruise ship can cover 17 nautical miles in 306 minutes. How many nautical miles will it travel in 234 minutes?
Answer by Cromlix(4381)

Hi there,
306 minutes to sail 17 nautical miles
1 minute to sail 17/306 = 1/18th of a nautical mile
234 minutes to sail 1/18 x 234 = 13 nautical miles.
Hope this helps :-)